OpenStudy (anonymous):

What are the first, second, and third frequencies audible from a 20 cm long organ pipe when (A) only one end is open and when (B) both ends are open. The speed of sound through the air inside the organ pipe is 343m/s. (Hint: if a harmonic does not exist, it would not be heard.)

OpenStudy (anonymous):

A) 1st harmonic: fn = (1*343)/(4*20) = 4.3 2nd harmonic does not exist because it even numbered harmonics do not exist when it is closed in one end and open in another. 3rd harmonic: fn = (3*343)/(4*20) = 17.15 B) 1st harmonic: fn = (1*343)/(2*20) = 8.6 2nd harmonic: fn = (2*343)/(2*20) = 17.2 3rd harmonic: fn = (3*343)/(2*20) = 25.7 is this correct?

OpenStudy (anonymous):

Looks good.